OpenStudy (anonymous):

Matt plans to put concrete on a rectangular portion of his driveway. The portion is 8 feet long and 4 inches high. The price of concrete is $98 per cubic yard. The total cost of the concrete Matt needs is $58.07. Which of the following is closest to the width of the portion of the driveway on which Matt plans to put concrete?

OpenStudy (anonymous):

Someone help please!

OpenStudy (ybarrap):

 How many feet in a yard? How many inches in a foot? How many cubic feet in a cubic yard? We'll need these to answer the question.

OpenStudy (ybarrap):



OpenStudy (anonymous):

This is the only thing that they gave me to answer it with [1 foot = 12 inches; 1 yard = 3 feet]

OpenStudy (ybarrap):

So, there are 3 feet in a yard: $$ \cfrac{3 \text{ feet}}{1 \text{ yard}} $$ So 9 cubic feet is 1 cubic yard: $$ \left (\cfrac{3 \text{ feet}}{1 \text{ yard}} \right )^3=\cfrac{9 \text{ feet}^3}{1 \text{ yard}^3} $$ Make sense so far?

OpenStudy (anonymous):

yes

OpenStudy (ybarrap):

The height of the concrete block is, since there is 12 inches in a foot: $$ 4 \text{ inches}\times\cfrac{1 \text{ foot}}{12 \text{ inches}}=\cfrac{4}{12} \text{ feet}\\ =\cfrac{1}{3}\text{feet} $$ Make sense?

OpenStudy (anonymous):

yupp.

OpenStudy (ybarrap):

k, so the volume of the concrete block is $$ w\times l\times h\\ w \text{ feet }\times 8 \text{ feet}\times \cfrac{1}{3}\text {feet}\\ =\cfrac{8w}{3} \text{ cubic feet} $$ But we need to convert to cubic yards, so since there are 9 cubic yards in a cubic foot: $$ =\cfrac{8w}{3} \text{ cubic feet}\times \cfrac{9 \text{ cubic yards}}{1 \text{ cubic feet}} \\=\cfrac{72w}{3}\text{ cubic yards} $$ Ok?

OpenStudy (ybarrap):

oops, that's 9 cubic feet in a cubic yard, so...

OpenStudy (anonymous):

Ok so what would you do next?

OpenStudy (ybarrap):

$$ =\cfrac{8w}{3} \text{ cubic feet}\times \cfrac{1 \text{ cubic yards}}{9 \text{ cubic feet}} \\=\cfrac{8w}{27}\text{ cubic yards} $$

OpenStudy (anonymous):

So you would divide 27 by 8 and get 3.375? But round it to 3 right?

OpenStudy (ybarrap):

We know the total cost is $58.07 and the cost per cubic yard is $98, so then $$ \cfrac{8w}{27}\text{ cubic yards}\times \cfrac{\text{\$}98}{1 \text{ cubic yard}}\\ =\cfrac{8\times98w}{27}=\text{\$}58.07\\ \implies w=\cfrac{58.07\times27}{8\times98} $$ You'll just need to calculate and round to the nearest hundredth -- I'll let you do that :-) Notice how all the units canceled at each step -- that's the key to solving all these problems. That's it!

OpenStudy (anonymous):

Thanks(:

OpenStudy (ybarrap):

yw